OpenStudy (anonymous):

The integral of 1/(x)^(-4/3) from 0 to 8?

OpenStudy (ipwnbunnies):

Rewrite the function so the x term is in the numerator.

OpenStudy (ipwnbunnies):

More like rewrite the integral.

OpenStudy (anonymous):

\[\int_0^8\frac{1}{x^{-4/3}}~dx=\int_0^8x^{4/3}~dx=\left[\frac{x^{4/3+1}}{\frac{4}{3}+1}\right]_0^8=\cdots\]